Samples of certain size are drawn from a normally distributed population with S. D. 16. It is observed that the probability

of the sample mean lying between 9.8 and 14.6 is 45.14%. Find the sample-size and the population mean. (Given that

∫ ( )


1
Expert's answer
2021-12-31T08:45:27-0500

The probability given is,

"p(9.8\lt\bar{x}\lt14.6)=0.4514", with "\sigma=16"

Since the sample mean is normally distributed, we standardize this probability as follows,

"p({(9.8-\mu)\over({16\over\sqrt{n}})}\lt Z\lt {(14.6-\mu)\over({16\over\sqrt{n}})})=0.4514"

Since the the two points given are symmetric about the population mean, it implies that to find probability for each point, we divide the given probability by 2. That is, "{0.4514\over 2}=0.2257".

Let, "{(9.8-\mu)\over({16\over\sqrt{n}})}" be "Z_0" and "{(14.6-\mu)\over({16\over\sqrt{n}})}" be "Z_1".

So,

"p(Z_0\lt Z\lt 0)=0.2257\implies\phi(0)-\phi(Z_0)=0.2257"

Since "\phi(0)=0.5,\phi(Z_0)=0.5-0.2257=0.2743"

From the standard normal tables, we can see that "Z_0=-0.6"

Also,

"p(0\lt Z\lt Z_1)=0.2257\implies \phi(Z_1)-\phi(0)=0.2257\implies \phi(Z_1)=0.2257+0.5=0.7257"

From the standard normal tables, we can see that "Z_1=0.6"

Therefore,

"Z_0={(9.8-\mu)\over({16\over\sqrt{n}})}=-0.6\implies 9.8\sqrt{n}-\mu\sqrt{n}=-9.6......(i)"

"Z_1= {(14.6-\mu)\over({16\over\sqrt{n}})}=0.6\implies14.6\sqrt{n}-\mu\sqrt{n}=9.6......(ii)"

Subtracting equation "(i)" from equation "(ii)" we have,

"4.8\sqrt{n}=19.2\implies\sqrt{n}={19.2\over4.8}\implies n=16"

We take any of the two equations and solve for "\mu",

Taking equation "(i),"

"9.8*4-4\mu=-9.6\implies (39.2+9.6)=4\mu\implies \mu=12.2"

Therefore, the population mean "\mu" and the sample size "n" are 12.2 and 16 respectively.